OpenStudy (cookielate):

Dale baby-sat and worked at a restaurant for a total of 11 hours. He made a total of $58. He earns $5 per hour babysitting and $6 per hour at the restaurant. How many hours did he work at each job? Variables x= # hours at restaurant y= # hours worked babysitting can someone be nice enough to help me with this question please anyone

OpenStudy (anonymous):

@mathstudent55

OpenStudy (cookielate):

thanks for asking others to help me

OpenStudy (anonymous):

np :)

OpenStudy (cookielate):

@mathstudent55

OpenStudy (cookielate):

@calculusfunctions

OpenStudy (cookielate):

@TheSmartOne

OpenStudy (cookielate):

@sleepyjess

OpenStudy (cookielate):

@AlexandervonHumboldt2

TheSmartOne (thesmartone):

We need to use the information they gave us, and make 2 equations and solve them.

OpenStudy (cookielate):

idk how to do elimination and system of equations

OpenStudy (cookielate):

ill try my best

OpenStudy (cookielate):

hours of restaurant is the total of 11 hours and gets total of $58

OpenStudy (cookielate):

@emcrazy14

OpenStudy (cookielate):

@dan815

TheSmartOne (thesmartone):

Key information: He earns $5 per hour babysitting and $6 per hour at the restaurant. Variables x= # hours at restaurant y= # hours worked babysitting Dale baby-sat and worked at a restaurant for a total of 11 hours.

TheSmartOne (thesmartone):

So one equation we can get from that information is, it says he works for a total of 11 hours. And it tells us that x is hours worked at a restaurant and y is hours worked babysitting. So we get the equation \(\sf x+y=11\)

TheSmartOne (thesmartone):

And it says that he makes $5 per hour for babysitting and $6 per hour for wokring in the restaraunt. And he gets a total of $58 So our second equation is \(\sf 6x+5y=58\)

TheSmartOne (thesmartone):

And solve those 2 equations. And whatever value you get for x is the number of hours he worked at the restaurant. And y is the value of how many hours he worked babysittingg

OpenStudy (cookielate):

so i will get two different answers??

OpenStudy (cookielate):

@perl

TheSmartOne (thesmartone):

no. You will get one answer for x. And one answer for y. You are supposed to get two answers.

OpenStudy (cookielate):

oh like subitute it

OpenStudy (cookielate):

he answer for this equation i think its x+y=-1 i think thats the answer -1

OpenStudy (cookielate):

**the

OpenStudy (cookielate):

hello

TheSmartOne (thesmartone):

Solve these two equations for x and y. \(\sf x+y=11\) \(\sf 6x+5y=58\)

TheSmartOne (thesmartone):

You can graph the 2 equation using Desmos, and find the point where they intersect which is (x,y) and your answer for x is the point (which is x) and for y is y. https://www.desmos.com/calculator

OpenStudy (cookielate):

so the answer is -1

OpenStudy (cookielate):

for this equation 6x+5y=58 when i graph it said 12 and 10 but idk which one it is

OpenStudy (cookielate):

hello

OpenStudy (cookielate):

@TheSmartOne hello??

OpenStudy (cookielate):

sorry

TheSmartOne (thesmartone):

At what point do they intersect at? http://prntscr.com/6ij7ow

OpenStudy (cookielate):

12 and 10

TheSmartOne (thesmartone):

Let's take a zoom in to that graph: http://prntscr.com/6ij86s

OpenStudy (cookielate):

oh nevermind its 3,8

TheSmartOne (thesmartone):

Exactly!

OpenStudy (cookielate):

so would it just only be 3

TheSmartOne (thesmartone):

And \(\sf (3,8)\) is in the form of \(\sf (x,y)\) So \(\sf x=3\\y=8\)

OpenStudy (cookielate):

so it would just be 3 for this equation 6x+5y= 58

OpenStudy (cookielate):

hello? @TheSmartOne

OpenStudy (cookielate):

hello

TheSmartOne (thesmartone):

The question wants two answers. It wants 2 answers. One for how many hours you worked at the restaurant (x) and one for how many hours you worked babysitting (y).

OpenStudy (cookielate):

so the two answers are -1 and 3 ?? im right or no

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @TheSmartOne And \(\sf (3,8)\) is in the form of \(\sf (x,y)\) So \(\sf x=3\\y=8\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (cookielate):

for the two equations ??

OpenStudy (cookielate):

or just one ?? last question

TheSmartOne (thesmartone):

For the 2 equations.

OpenStudy (cookielate):

oh ok thanks sorry for bothering you and thats the solution right x=3 and y=8 right thats the solution thanks alot for helping me

TheSmartOne (thesmartone):

Np :)